# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v4

@@ 3 QRNG based on Single-photon Source
3.1 Effect of Imperfect Devices
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section 5. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
@@ 3.1 Effect of Imperfect Devices
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section 8 for a discussion on these.
4 Existing QRNG Models based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging [54, 55]. However, different experimental approaches [56, 57, 58, 59, 60, 61, 62, 63] have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval [96]. In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
@@ 4 Existing QRNG Models based on Coherent Pulse
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible [96]. A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
@@
But in this process [13], the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature [12, 15, 14, 16, 17, 17]. In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
@@
Therefore, the random number generated with this procedure follows exponential distribution [15, 14, 16, 17]. Another QRNG scheme is proposed in [12], as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
5 Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section 2 we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@ 5 Our Proposal of Two-fold Statistical Method for Testing Quantumness
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. The statistical tools have been used in this section is discussed in Appendix C.
5.1 First Phase: Interval Estimation
@@
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$  [97], where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
@@
is an unbiased and consistent estimate of $\sigma^{2}$ [97]. Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
@@
5.2 Hypothesis Testing as an Alternative to Interval Estimation
@@
5.3 Second Phase: Goodness-of-fit Test
@@
Figure 4: Probability density function for $\chi^{2}$ distribution. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by stars, and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test [97, 98] on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
@@
is the well-known Gamma function [97]. The cumulative distribution function in this case is given by,
@@
6 Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (12) of Section 4 consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section 7.
6.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
Currently available photon arrival time-based QRNGs [37, 38] use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ 6.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
This was also observed by Nie et al. [17, Eq. (2)]. However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
6.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Until now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, the first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ 6.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
is the probability of detecting no photon in a complete reference cycle. In that case, this reference cycle has no contribution to the random number. Therefore the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins. Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Note that $f(1)$ is the probability $P_{1}$ in Ref. [17, Eq. (4)]. As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section 7. Figure 5 shows how $\mu Td$ affects the probability (31). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section 7.
@@
Since in the probability distribution of Eq. (6.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (34) is identical with Eq. (6.2). Therefore, the probability mass function corresponding to the probability Eq. (34) is also given by Eq. (31) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (12)). Since the distributions given by Eq. (31) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection 5.3 to test the quantumness in the raw data generated by a QRNG.
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution given by Eq. (31) is $\epsilon$-random. Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
Definition 1
@@
Therefore, the distribution (31) becomes $\epsilon$-random if and only if
@@
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (31). These test results are provided in Section 8.
7 Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (32) of Subsection 6.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
@@ 7 Effect of $\mu,T,d$ on random number generation rate and cost
The Shannon entropy [99, 100] of a uniform distribution over $n$ points is given by
@@
A detailed derivation of this entropy along with the definition of Shannon entropy and minimum entropy has been added in Appendix D. Clearly, $H(Q)\neq\log N=H(U)$. So, we have to apply post-processing as mentioned in Eq. (37) to achieve $H(E)=H(U)$ for some uniform distribution $U$ whose support contains less than $N$ points. This post-processing, also known as randomness extractor, is a data compression procedure using universal hashing [101]. The minimum entropy $H_{min}$ of the raw data decides the amount of this data compression. If $H_{min}\geq\frac{m}{n}$ for some positive integers $m,n$, an $n$-bit raw data needs to be compressed to an $m$-bit data to make it uniform random [17]. As $f(i)$ in Eq. (31) is a decreasing function, the minimum entropy of the photon arrival-time-based QRNG discussed above would be
$H_{min}$ $=-\log_{2}f(1)$
$=-\log_{2}\frac{1-e^{-\frac{\mu Td}{N}}}{1-e^{-\mu Td}}$
$=-\log_{2}\left(1-e^{-\frac{\mu Td}{N}}\right)+\log_{2}\left(1-e^{-\mu Td}\right).$ (41)
Nie et al. [17, Eq. (5)] gave a lower bound of this $H_{min}$. Since our aim is to reduce the amount of post-processing, which depends on $\mu Td$, we want to investigate $\mu,T$ and $d$ separately.
@@
7.1 Effect of the Expected Photon Numbers

### v5

@@ III QRNG based on Single-photon Source
III.1 Effect of Imperfect Devices
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section V. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
@@ III.1 Effect of Imperfect Devices
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section IX for a discussion on these.
IV Existing QRNG Models based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging PhysRevLett.72.210 ; kim1999single . However, different experimental approaches PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval argillander2023quantum . In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
@@ IV Existing QRNG Models based on Coherent Pulse
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible argillander2023quantum . A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
@@
But in this process 10.1063/1.2961000 , the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature 10.1063/1.2720728 ; WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
@@
Therefore, the random number generated with this procedure follows exponential distribution WayneQRNG ; WayneRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 . Another QRNG scheme is proposed in 10.1063/1.2720728 , as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
V Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section II we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@ V Our Proposal of Two-fold Statistical Method for Testing Quantumness
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. Numerous photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and experimentally demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years, providing viable tools for carrying out these tests. The statistical tools have been used in this section is discussed in C.
V.1 First Phase: Interval Estimation
@@
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$ hogg2019introduction , where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
@@
is an unbiased and consistent estimate of $\sigma^{2}$ hogg2019introduction . Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
@@
V.2 Hypothesis Testing as an Alternative to Interval Estimation
@@
V.3 Second Phase: Goodness-of-fit Test
@@
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test hogg2019introduction ; chisq on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
@@
is the well-known Gamma function hogg2019introduction . The cumulative distribution function in this case is given by,
@@
Figure 4: Probability density function (22) for $\chi^{2}$ distribution. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by stars, and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
@@
VI Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (12) of Section IV consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section VII.
VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
Currently available photon arrival time-based QRNGs quRNG ; TROPOS use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
This was also observed by Nie et al. (10.1063/1.4863224, , Eq. (2)). However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Til now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, the first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
is the probability of detecting no photon in a complete reference cycle. In that case, this particular reference cycle has no contribution to the random number. Therefore, the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins given by Eq. (31). Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Note that $f(1)$ is the probability $P_{1}$ in Ref. (10.1063/1.4863224, , Eq. (4)). As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section VII. Figure 5 shows how $\mu Td$ affects the probability (31). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section VII.
@@
Since in the probability distribution of Eq. (VI.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (34) is identical with Eq. (VI.2). Therefore, the probability mass function corresponding to the probability Eq. (34) is also given by Eq. (31) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (12)). Since the distributions given by Eq. (31) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection V.3 to test the quantumness in the raw data generated by a QRNG.
Definition 1.
@@
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution is $\epsilon$-random given by Eq. (36). Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
The distribution (31) becomes $\epsilon$-random if and only if
@@
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (31). These test results are provided in Section IX.
VII Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (32) of Subsection VI.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
@@ VII Effect of $\mu,T,d$ on random number generation rate and cost
The Shannon entropy 6773024 ; 6773067 of a uniform distribution over $n$ points is given by
@@
A detailed derivation of this entropy along with the definition of Shannon entropy and minimum entropy has been added in D. Clearly, $H(Q)\neq\log N=H(U)$. So, we have to apply post-processing as mentioned in Eq. (37) to achieve $H(E)=H(U)$ for some uniform distribution $U$ whose support contains less than $N$ points. This post-processing, also known as randomness extractor, is a data compression procedure using universal hashing PhysRevA.87.062327 . The minimum entropy $H_{min}$ of the raw data decides the amount of this data compression. If $H_{min}\geq\frac{m}{n}$ for some positive integers $m,n$, an $n$-bit raw data needs to be compressed to an $m$-bit data to make it uniform random 10.1063/1.4863224 . As $f(i)$ in Eq. (31) is a decreasing function, the minimum entropy of the photon arrival-time-based QRNG discussed above would be
$H_{min}$ $=-\log f(1)=-\log\frac{1-e^{-\frac{\mu Td}{N}}}{1-e^{-\mu Td}}$
$=-\log\left(1-e^{-\frac{\mu Td}{N}}\right)+\log\left(1-e^{-\mu Td}\right).$ (41)
Nie et al. (10.1063/1.4863224, , Eq. (5)) gave a lower bound of this $H_{min}$. Since our aim is to reduce the amount of post-processing, which depends on $\mu Td$, we want to investigate $\mu,T$ and $d$ separately.
@@
VII.1 Effect of the Expected Photon Numbers